OpenStudy (anonymous):

Predict the next two numbers in this sequence: 1/4, 1/9, 1/16, 1/25,...

OpenStudy (turingtest):

Seems like the denominator increases by 2 more each time. Looking just at the denominator: 4 9=4+5 16=9+7 25=16+9 so the pattern is +5,+7,+9... to the denominator. Take it from there.

OpenStudy (anonymous):

Ok. Thanks alot!